# CleanPose: Category-Level Object Pose Estimation via Causal Learning and Knowledge Distillation

### v1

@@ Abstract
Category-level object pose estimation aims to recover the rotation, translation and size of unseen instances within predefined categories. In this task, deep neural network-based methods have demonstrated remarkable performance. However, previous studies show they suffer from spurious correlations raised by “unclean” confounders in models, hindering their performance on novel instances with significant variations. To address this issue, we propose CleanPose, a novel approach integrating causal learning and knowledge distillation to enhance category-level pose estimation. To mitigate the negative effect of unobserved confounders, we develop a causal inference module based on front-door adjustment, which promotes unbiased estimation by reducing potential spurious correlations. Additionally, to further improve generalization ability, we devise a residual-based knowledge distillation method that has proven effective in providing comprehensive category information guidance. Extensive experiments across multiple benchmarks (REAL275, CAMERA25 and HouseCat6D) hightlight the superiority of proposed CleanPose over state-of-the-art methods. Code will be released.
@@ 1 Introduction
Figure 1: Comparison of (a) existing pose estimation approaches with (b) the proposed causal learning pipeline. The front-door adjustment is used to mitigate the negative effect from confounders. We also leverage rich 3D semantic knowledge to provide comprehensive category guidance.
Category-level object pose estimation (COPE) aims to predict the 3D rotation, 3D translation and 3D metric size, for arbitrary objects within predefined categories. This task, unlike instance-level pose estimation [47, 64, 12, 9, 19, 24], dose not require high-quality CAD models, making it feasible to perceive a broader range of novel objects instead of a single instance. In recent years, this task has garnered significant attention due to its essential role in various practical applications, including augmented reality [34], autonomous driving [3] and robotic manipulation [44], etc.
The core challenge of COPE lies in the complicated and diverse intra-category variations. To obtain robust category features and achieve accurate pose prediction, previous methods [5, 22, 62, 23, 26, 29, 25, 4, 63] largely rely on deep neural networks (DNNs) due to their significant learning and modeling capabilities, offering impressive progress. Although DNNs have demonstrated extraordinary performance in COPE, recent studies [16, 49] reveal that they may focus on spurious correlations to benefit predictions, as their optimization objective is to learn statistical distribution only. These illogical correlations in models are typically raised by the underlying confounders [38] behind the task, e.g., dataset biases [17, 60]. Here, confounders are variables that influence both inputs and outcomes. DNN-based pose estimation models are sensitive to the confounders during data fitting, as shown in Fig. 1(a). For instance, the COPE models may overfit to specific objects’ appearance and poses due to the lack of data variety in dataset, damaging the generalization to novel instances
Another important aspect of COPE neglected by previous works is the comprehensive category-specific guidance. Most of them [5, 23, 28, 25, 4] solely derive category information from a limited instances per category in training set, which may also capture only limited category diversity. One way to tackle above two issues is to create broader and more diverse datasets [7, 50, 17, 60], though valuable, achieving a perfectly balanced dataset free of bias remains nearly impossible. Additionally, the dataset’s scale is still significantly constrained by the cost of 3D data annotation [61].
@@
To better address the dilemma, we direct our attention to the human observation mechanism. In fact, the reason why humans can effectively handle variations among intra-category objects is that we learn inherent causality beyond biased observation, achieving excellent analogical association capability. Motivated by this finding, for the first time, we propose to incorporate the causal inference [36, 37] into the formulation of COPE. In this way, we can investigate the causal relation among variables and equip COPE models with similar cognitive abilities that human have. Technically, it is non-trivial to incorporate the causal learning in COPE applications because of the following challenges: (i) The unique modality of 3D data makes it impractical to directly apply existing causal modeling techniques [15]. (ii) Subsequently, the confounders in pose estimation task are inherently unobserved and elusive, which further increases the challenge of identifying these confounders and mitigating their negative impacts. (iii) Moreover, enhancing the model’s category generalization capability can not be overlooked. How to improve the generalization while simultaneously deconfoundering needs to be carefully considered.
@@
To this end, we present CleanPose, a concise yet effective framework with causal learning [36] and knowledge distillation to enhance the category-level pose estimation, as shown in Fig. 1(b). Faced with the challenge that hidden confounders are elusive or even unobserved in our task, we propose to develop a causal inference approach based on front-door adjustment, which can effectively approximate the predicted expectations by potential confounders. Moreover, to create suitable dictionary for representing confounders, we devise and maintain the dictionary as a dynamic queue, similar with MoCo [11], to efficiently update training samples. To further address the weakness in category information guidance, we design a residual-based feature knowledge distillation layer to transfer abundant point cloud category information of powerful 3D foundation model, ULIP-2 [53], to guide the category feature learning within the model. As demonstrated by extensive experiments, our findings reveal the impact of integrating causal learning to deconfound biases, and providing comprehensive category information on COPE, enhancing the model’s robustness and generalization.
@@
  • We propose CleanPose, a pioneering solution to mitigate the confoundering effect in category-level pose estimation via causal learning. Taking inspiration from human observation mechanism, we propose to identify the causal effect to achieve unbiased estimation, recovering correct pose of novel instances within category.
  • We develop a residual category knowledge distillation approach to transfer rich 3D category knowledge from 3D foundation models into category-level networks, enhancing the intra-category generalization.
  • Our proposed CleanPose achieves state-of-the-art performance on three mainstream challenge benchmarks (REAL275 [48], CAMERA25 [48] and HouseCat6D [17]). For instance, the accuracy attains 61.5% on rigorous metric 5°2 cm of REAL275 dataset, surpassing the current best method with a large margin by 4.5%.
@@ 2 Related Works
Category-level Object Pose Estimation. The objective of this task encompasses predicting the 9DoF pose for unseen objects within predefined categories. To address this challenging task, pioneer method NOCS [48] suggests mapping input shape to a normalized canonical space and recovering the pose via Umeyama algorithm [45]. SPD [43] proposes a method for deriving and utilizing the shape prior for each category. This crucial insight inspires many subsequent prior-based works [20, 6], which further improve the use of shape priors, continuously improving the pose estimation performance. More recently, prior-free methods [5, 62, 23, 4, 25] have achieved impressive performance. VI-Net [23] separates rotation into viewpoint and in-plane rotations, while SecondPose [4] propose to extract hierarchical panel-based geometric features for point cloud. AG-Pose [25] achieves current state-of-the-art performance by explicitly extract local and global geometric keypoint information of different instances. However, above methods have not fully considered that DNNs lack the ability to perform causal inference, which may leads the models to learn spurious correlations. In addition, most of them either explicitly or implicitly learn category information from training set, which results in poor generalization ability to unseen instances within the categories.
@@
Knowledge Distillation. Knowledge distillation [13] is a technique that transfers knowledge from a teacher model to a student model. Transferring the knowledge of foundation models for downstream tasks has been proven to be effective [8, 14, 66]. ViLD [8] distill knowledge from CLIP [41] to achieve open-vocabulary object detection, while some methods [14, 66] utilize the generalizability of VLMs to address the video recognition task. The most related method to our work is CLIPose [26]. It aligns the representations of the image, point cloud, and text modalities through multi-modal contrastive learning. However, such semantic labels cannot effectively guide the perception of point cloud structural information. Unlike CLIPose, our method distills knowledge directly from the point cloud encoder of 3D foundation models via a residual MLP layer, providing more comprehensive guidance for category learning.
@@
3 Preliminary
@@ 3.1 Task Formulation
Given an RGB-D image containing objects from a predefined set of categories, off-the-shelf segmentation models such as MaskRCNN [10] are employed to obtain masks and category labels for each object in the images. Then, the segmentation masks can be utilized to get the cropped RGB image $\mathcal{I}_{obj}\in\mathbb{R}^{H\times W\times 3}$ and the point cloud $\mathcal{P}_{obj}\in\mathbb{R}^{N\times 3}$, where $N$ is the number of points and $\mathcal{P}_{obj}$ is acquired by back-projecting the cropped depth image with camera intrinsics followed by a downsampling process. With the input $\mathcal{I}_{obj}$ and $\mathcal{P}_{obj}$, the objective of COPE [48] is to recover the 9DoF poses, including the 3D rotation $\mathcal{R}\in SO(3)$, the 3D translation $t\in\mathbb{R}^{3}$ and 3D metric size $s\in\mathbb{R}^{3}$.
@@ 3.2 The Causal Modeling of CleanPose
To quantify the underlying logic behind human observation, we construct a structural causal model [36, 37] capturing the relationships among the key variables in COPE: visual input $\mathcal{X}$, output pose $\mathcal{Y}$, mediator $\mathcal{M}$ and hidden confounders $\mathcal{U}$. We illustrate the causal model in Fig. 2, where each direct link denotes a causal relationship between two nodes.
@@
  • $\mathcal{X}\leftarrow\mathcal{U}\rightarrow\mathcal{Y}$ (Hidden confounders): The confounders are extraneous variables that influence both inputs and outputs, e.g., dataset biases [17, 60] or category specific attributes [43]. $\mathcal{U}\rightarrow\mathcal{X}$ exists because the input data is inevitably affected by the limited resources in real world and sampling noises from sensors when collection and simulation. Moreover, $\mathcal{U}\rightarrow\mathcal{Y}$ emerges because collected scenes, annotation bias, or the variety of pose also affect the probability of pose distributions. Traditional DNNs-based COPE methods tend to model the statistical correlations $P(\mathcal{Y}|\mathcal{X})$, given the optimization goal of maximizing the pose accuracy [63]. With the lack of modeling of hidden confounders, no matter how large the amount of training data is, the model can not identify the ture causal effect from $\mathcal{X}$ to $\mathcal{Y}$.
@@ 4.1 Feature Extractor
Following [25], we utilize the PointNet++ [40] to extract point feature $\mathcal{F}_{P}\in\mathbb{R}^{N\times C_{1}}$ of input point cloud $\mathcal{P}_{obj}$. As for the RGB image $\mathcal{I}_{obj}$, we adopt DINOv2 [35] as our image feature extractor, which has been proven to extract abundant semantic-aware information from RGB images [39]. We select those pixel features corresponding to $\mathcal{P}_{obj}$ and utilize linear interpolation to propagate the original DINOv2 features into the final RGB features $\mathcal{F}_{I}\in\mathbb{R}^{N\times C_{2}}$. Eventually, we concatenate $\mathcal{F}_{P}$ and $\mathcal{F}_{I}$ to form $\mathcal{F}_{obj}\in\mathbb{R}^{N\times C}$ as the input for the subsequent networks.
@@ 4.2 Causal Inference of CleanPose
The Adjustment Formulation. To avoid making the pose estimation process overfit to specific hidden confounders, we propose to perform robust modeling according to the causal graph and exploit intervention [36] to alleviate the negative effects raised by confounders. Specifically, we employs do-operation [37], which provides scientifically sound methods for determining causal effects, as shown in Fig. 3(b). The specific process of intervention via do-operation is as follow. Based on Bayes’s theorem, the typical observational likelihood is as:
@@
For convenience, we define the network module as a linear function $f(x,\bm{x^{\prime}})$ to model the computation of $x$ and $\bm{x^{\prime}}$ to obtain $\mathcal{Y}$. Meanwhile, according to the Bayes’ rule and the the definition of Expected values, the front-door adjustment $\mathcal{FD}$(·) of Eq. 7 can be expressed as:
@@
To this end, we draw inspiration from MoCo [11] and devise a features sampling approach based on a dynamic queue. Specifically, we first utilize the 3D encoders of ULIP-2 [53] to extract a specified number of features for each category, constructing an initial 2D queue with a shape of $N_{c}\times N_{q}$, where $N_{c}$ and $N_{q}$ represent the number of categories and the queue length of each category, respectively. Next, we randomly sample $N_{s}$ features from the queue to form $\bm{x^{\prime}}$, which is denoted as $\mathcal{F}_{samp}\in\mathbb{R}^{N_{s}\times C}$ in the networks. Based on the queue’s characteristics, we update the keypoint features $\mathcal{F}_{kpt}$ from the current epoch into the queue via first-in-first-out (FIFO) approach, ensuring feature consistency within the queue. Following Eq. 10 to Eq. 12, the causality-enhanced features $\mathcal{F}_{f}$ are obtained as follow:
@@
4.3 Category Knowledge Distillation
As mentioned in above section, implicitly learning category knowledge from a limited dataset only results in limited intra-category generalization. Inspired by the feature alignment design [14, 66, 65] to transfer visual-language knowledge to different tasks, we propose a modified residual approach to distill the category knowledge from 3D foundation models into COPE network, as shown in Fig. 3(c). Unlike CLIPose [26] exploits contrastive learning to extract semantic knowledge from text and image modalities, we leverage the pre-trained 3D encoder in ULIP-2 [53], which possesses strong awareness of point cloud structures and category information. Specifically, given the input point cloud of a object $\mathcal{P}_{obj}$, we can directly feed it into the frozen 3D Encoder $\Phi_{ULIP}$ of ULIP-2, which can be written as:

### v2

@@ Abstract
In the effort to achieve robust and generalizable category-level object pose estimation, recent methods primarily focus on learning fundamental representations from data. However, the inherent biases within the data are often overlooked: the repeated training samples and similar environments may mislead the models to over-rely on specific patterns, hindering models’ performance on novel instances. In this paper, we present CleanPose, a novel method that mitigates the data biases to enhance category-level pose estimation by integrating causal learning and knowledge distillation. By incorporating key causal variables (structural information and hidden confounders) into causal modeling, we propose the causal inference module based on front-door adjustment, which promotes unbiased estimation by reducing potential spurious correlations. Additionally, to further confront the data bias at the feature level, we devise a residual-based knowledge distillation approach to transfer unbiased semantic knowledge from 3D foundation model, providing comprehensive causal supervision. Extensive experiments across multiple benchmarks (REAL275, CAMERA25 and HouseCat6D) hightlight the superiority of proposed CleanPose over state-of-the-art methods. Code will be available at https://github.com/chrislin0621/CleanPose.
@@ 1 Introduction
Figure 1: Comparison of (a) existing pose estimation approaches and (b) the proposed causal learning pipeline. The causal inference is employed to mitigate the negative effect raised by bias, and 3D semantic knowledge distilled from foundation model is leveraged to provide comprehensive causal guidance.
Category-level object pose estimation (COPE) [42] aims to predict the 9Dof pose for arbitrary objects within predefined categories. This task, unlike instance-level pose estimation [41, 58, 11, 8, 21], does not require high-quality CAD models, making it feasible to perceive a broader range of novel objects instead of a single instance. The complex intra-category variations have long been regarded as main challenge of COPE, thus recent methods [1, 17, 4, 19, 56, 20, 23, 22, 3] mainly focus on model designing to capture robust category features from data. Despite remarkable development, the performance of COPE models remains limited [23, 20], suggesting that feature learning may not be the fundamental issue. This prompts us to identify the key hindrance blocking further progress. We observe that inherent biases are pervasive in current datasets [42], e.g., repeated training samples and limited pose variety, which may mislead COPE models to overfit to familiar object’s appearance and poses during data fitting, as shown in Fig. 1(a). However, existing efforts often overlook these underlying data biases and the resulting negative impacts.
One way to tackle bias issue is to create broader and more diverse datasets [6, 44, 15, 54]. Though valuable, achieving a perfectly balanced dataset free of bias remains nearly impossible. Additionally, the dataset’s scale is still significantly constrained by the cost of 3D data annotation [55]. Therefore, the extension of datasets does not fundamentally solve the hindrance, developing a causal COPE models that can effectively confront and alleviate biases becomes a primary challenge.
To address this issue, we direct our attention to the human observation mechanism. In fact, the reason why humans can effectively handle variations among intra-category objects is that we can leverage the analogical association to infer the structural features and underlying causal relations. Moreover, the observations of numerous objects enable humans to maintain a stable perception of novel objects across different environments and viewpoints.
@@
Motivated by this finding, for the first time, we propose to incorporate the causal inference [32] into the formulation of COPE, exploiting the concepts of intervention [33] to mitigate the negative effects caused by confounders. Here, confounders are variables that influence both input and outcomes, just as dataset biases in this task simultaneously affect input sampling and output pose. The causal learning can investigate the causal relation among variables and equip COPE models with similar cognitive abilities that humans have. However, it is non-trivial to directly incorporate the causal inference in COPE applications because of the following challenges: (i) The unique modality of 3D data makes it impractical to directly apply existing causal modeling techniques [14]. (ii) Subsequently, the confounders in pose estimation task are inherently unobserved and elusive, which further increases the challenge of identifying these confounders. (iii) Moreover, limited training samples raise the difficulties in learning comprehensive causal relations.
@@
To overcome above challenges, we present CleanPose, a concise yet effective framework with causal learning and knowledge distillation to enhance the category-level pose estimation, as shown in Fig. 1(b). Faced with the challenge that hidden confounders are elusive or even unobserved in our task, we have two key designs: (i) The first design is a causal inference module based on front-door adjustment, which can effectively approximate the predicted expectations by potential confounders. Moreover, to suitably represent confounders, we devise and maintain a dynamic queue to efficiently update training samples, similar with MoCo [10]. To further mitigate biases, (ii) our second design is a residual-based feature knowledge distillation module, which can transfer unbiased point cloud category information of the 3D foundation model, ULIP-2 [47], into our model. Since the foundation model is pretrained on a diverse range of objects, it captures robust semantic features that can implicitly provide comprehensive causal supervision at the feature level. As demonstrated by extensive experiments, our findings reveal the impact of integrating causal learning to reduce biases, and providing comprehensive causal supervision, enhancing the model’s robustness and generalization.
@@
  • We propose CleanPose, the first solution to mitigate the confoundering effect in category-level pose estimation via causal learning. Taking inspiration from human observation mechanism, we propose to identify the causal effect to achieve unbiased estimation, recovering correct pose of novel instances within predefined categories.
  • To provide comprehensive causal supervision and enhance intra-category generalization, we develop a residual category knowledge distillation approach to transfer rich and unbiased 3D semantic knowledge from 3D foundation models into COPE networks.
  • Our proposed CleanPose achieves state-of-the-art performance on three mainstream challenge benchmarks (REAL275 [42], CAMERA25 [42] and HouseCat6D [15]). For instance, the accuracy attains 61.7% on rigorous metric 5°2 cm of REAL275 dataset, surpassing the current best method with a large margin by 4.7%.
@@ 2 Related Works
Category-level Object Pose Estimation. The objective of this task encompasses predicting the 9DoF pose for unseen objects within predefined categories. To address this challenging task, pioneer method NOCS [42] suggests mapping input shape to a normalized canonical space and recovering the pose via Umeyama algorithm [39]. SPD [38] proposes a method for deriving and utilizing the shape prior for each category. This crucial insight inspires many subsequent prior-based works [18, 5, 17], which further improve the use of shape priors, continuously improving the pose estimation performance. More recently, prior-free methods [4, 56, 20, 3, 22] have achieved impressive performance. VI-Net [20] separates rotation into viewpoint and in-plane rotations, while SecondPose [3] propose to extract hierarchical panel-based geometric features for point cloud. AG-Pose [22] achieves current state-of-the-art performance by explicitly extract local and global geometric keypoint information of different instances. However, above methods have not fully considered the dataset biases behind the task, which may leads the models to learn spurious correlations, damaging the generalization ability to unseen instances within the categories.
@@
Knowledge Distillation. Knowledge distillation [12] is a technique that transfers knowledge from a teacher model to a student model. Transferring the knowledge of foundation models for downstream tasks has been proven to be effective [7, 13, 60]. ViLD [7] distill knowledge from CLIP [36] to achieve open-vocabulary object detection, while some methods [13, 60] utilize the generalizability of VLMs to address the video recognition task. In this work, we leverage the 3D foundation model ULIP-2 [47] as a teacher model to guide the student model in correcting biases at the feature level, enhancing the robustness and generalization capabilities of COPE model from multiple perspectives.
3 Preliminaries
@@ 3.1 Task Formulation
Given an RGB-D image containing objects from a predefined set of categories, off-the-shelf segmentation models such as MaskRCNN [9] are employed to obtain masks and category labels for each object in a image. Then, the segmentation masks can be utilized to get the cropped RGB image $\mathcal{I}_{obj}\in\mathbb{R}^{H\times W\times 3}$ and the point cloud $\mathcal{P}_{obj}\in\mathbb{R}^{N\times 3}$, where $N$ is the number of points and $\mathcal{P}_{obj}$ is acquired by back-projecting the cropped depth image with camera intrinsics followed by a downsampling process. With the input $\mathcal{I}_{obj}$ and $\mathcal{P}_{obj}$, the objective of COPE [42] is to recover the 9DoF poses, including the 3D rotation $\mathcal{R}\in SO(3)$, the 3D translation $t\in\mathbb{R}^{3}$, and 3D metric size $s\in\mathbb{R}^{3}$.
@@ 3.2 The Causal Modeling of CleanPose
To quantify the underlying logic behind human observation, we construct a structural causal model [32, 33] capturing the relationship among the key variables in COPE: visual input $\mathcal{X}$, output pose $\mathcal{Y}$, mediator $\mathcal{M}$ and hidden confounders $\mathcal{U}$. We illustrate the causal model in Fig. 2, where each direct link denotes a causal relationship between two nodes.
@@
  • $\mathcal{X}\leftarrow\mathcal{U}\rightarrow\mathcal{Y}$ (Hidden confounders): The confounders are extraneous variables that influence both inputs and outputs, e.g., dataset biases [15, 54] or category specific attributes [38]. $\mathcal{U}\rightarrow\mathcal{X}$ exists because the input data is inevitably affected by the limited resources in real world and sampling noises from sensors when collection and simulation. Moreover, $\mathcal{U}\rightarrow\mathcal{Y}$ emerges because collected scenes, annotation bias, or the variety of pose also affect the probability of pose distributions. Existing COPE methods tend to model the statistical correlations $P(\mathcal{Y}|\mathcal{X})$, given the optimization goal of maximizing the pose accuracy [57]. With the lack of modeling of hidden confounders, no matter how large the amount of training data is, the model can not identify the true causal effect from $\mathcal{X}$ to $\mathcal{Y}$.
@@ 4.1 Feature Extractor
Following [22], we utilize the PointNet++ [35] to extract point feature $\mathcal{F}_{P}\in\mathbb{R}^{N\times C_{1}}$ of input point cloud $\mathcal{P}_{obj}$. As for the RGB image $\mathcal{I}_{obj}$, we adopt DINOv2 (ViT-S/14) [31] as our image feature extractor, which has been proven to extract abundant semantic-aware information from RGB images [34]. We select those pixel features corresponding to $\mathcal{P}_{obj}$ and utilize linear interpolation to propagate the original DINOv2 features into the final RGB features $\mathcal{F}_{I}\in\mathbb{R}^{N\times C_{2}}$. Eventually, we concatenate $\mathcal{F}_{P}$ and $\mathcal{F}_{I}$ to form $\mathcal{F}_{obj}\in\mathbb{R}^{N\times C}$ as the input for the subsequent networks.
@@ 4.2 Causal Inference of CleanPose
The Adjustment Formulation. To avoid making the pose estimation process overfit to specific hidden confounders, we propose to perform causal modeling according to the causal graph and the concepts of intervention [32]. Specifically, the process of intervention can be reflected as do-operation [33], which provides scientifically sound methods for determining causal effects, as shown in Fig. 3(b). The specific casual modeling via do-operation is as follow. Based on Bayes’s theorem, the typical observational likelihood is as:
@@
For convenience, we define the network module as a linear function $f(x,\bm{x^{\prime}})$ to model the computation of $x$ and $\bm{x^{\prime}}$ to obtain $\mathcal{Y}$. Meanwhile, according to the Bayes’ rule and the definition of Expected values, the front-door adjustment $\mathcal{FD}$(·) of Eq. 7 can be expressed as:
@@
To this end, we draw inspiration from MoCo [10] and devise a features sampling approach based on a dynamic queue. Specifically, we first construct a 2D queue with a shape of $N_{c}\times N_{q}$, where $N_{c}$ and $N_{q}$ represent the number of categories and the queue length of each category, respectively. We then select training instances at random equal to $N_{q}$ for each category, extract their features using 3D encoders of ULIP-2 [47], and store them into the queue to complete the initialization. Subsequently, we randomly sample $N_{s}$ features from the queue to form $\bm{x^{\prime}}$, which is denoted as $\mathcal{F}_{samp}\in\mathbb{R}^{N_{s}\times C}$ in the networks. To maintain feature consistency within the queue, we update it using the output features $\mathcal{F}_{f}$ from the current epoch using a first-in-first-out (FIFO) strategy. Following Eq. 10 to Eq. 12, the causality-enhanced features $\mathcal{F}_{f}$ are obtained as follow:
@@
4.3 Knowledge Distillation
As mentioned in above section, the foundation model is trained on a broader dataset covering a wide range of categories and scenarios, enabling it to be exposed to diverse data distributions and to implicitly learn more robust debiased feature representations, which can supervise the COPE model to confront biases at the feature level. Inspired by the feature alignment design [13, 60, 59] to transfer visual-language knowledge to different tasks, we exploit 3D foundation model ULIP-2 [47] as the teacher model to provide causal guidance for COPE model. Specifically, we introduce a residual-based approach to distill debiasing knowledge from 3D foundation models into COPE network, as shown in Fig. 3(c). Unlike CLIPose [23] exploits contrastive learning to extract semantic knowledge from text and image modalities, we leverage the pre-trained 3D encoder of ULIP-2, which possesses strong awareness of point cloud structures and category information. Specifically, given the input point cloud of a object $\mathcal{P}_{obj}$, we can directly feed it into the frozen 3D Encoder $\Phi_{ULIP}$ of ULIP-2, which can be written as: